OpenStudy (gabylovesyou):

Identify the 16th term of a geometric sequence where a1 = 4 and a8 = −8,748. −172,186,884 −57,395,628 57,395,628 172,186,884

OpenStudy (anonymous):

You still post questions here? smh

OpenStudy (gabylovesyou):

@TheInspector yep cx

OpenStudy (gabylovesyou):

@johnweldon1993

OpenStudy (johnweldon1993):

Any ideas? Knowing the formula we want to use is \(\large a_n = a_1 * r^{n - 1}\)

OpenStudy (gabylovesyou):

4 * ??

OpenStudy (shadowlegendx):

\[a^8 / a^1 = ar^7 / a = r^7\]

OpenStudy (shadowlegendx):

Youʻre on the right track

OpenStudy (johnweldon1993):

First we need to solve for the 'r' *common ratio* So plug in what we know \[\large a_8 = a_1 * r^{8-1}\] \[\large -8748 = 4r^{7}\] Since we knew what a8 and a1 were...what would solving that for 'r' result in?

OpenStudy (gabylovesyou):

;c idk. im so lost.

OpenStudy (shadowlegendx):

Common Ratio: 7th root of \[\frac{ -8748 }{ 4 } = -3\]

OpenStudy (johnweldon1993):

Alright, so starting from the beginning...to find the n'th term in a geometric sequence...you use the formula \(\large a_n = a_1 * r^{n-1}\) This says...to find the n'th term...take the first term in the sequence and multiply it by the common ratio raised to the n'th - 1 power So...we are given \(\large a_8 = -8748\) and \(\large a_1 = 4\) So we need to solve for 'r' first Plug in what we know \[\large -8748 = 4r^{8-1}\] which becomes \[\large -8748 = 4r^7\] Or \[\large r = \sqrt[7]{-8748}\] What does r equal?

OpenStudy (johnweldon1993):

oops...my mistake up there...i never divided by the 4!

OpenStudy (johnweldon1993):

\[\large r = \sqrt[7]{\frac{-8748}{4}}\]

OpenStudy (shadowlegendx):

haha, yeah

OpenStudy (gabylovesyou):

idk how to solve for r tho..

OpenStudy (gabylovesyou):

-8748 = 4r^7 i understand how we got this.. now what

OpenStudy (johnweldon1993):

Okay so from that step...divide both sides by 4 \[\large -2187 = r^7\] So to solve for 'r'...we take the 7th root of that number

OpenStudy (johnweldon1993):

Same thing as taking a square root basically we know \(\large \sqrt[2]{4} = 2\) This time it's just a bigger number...but just plug it into your calculator :)

OpenStudy (johnweldon1993):

Or google of course :P and you'll see it comes out to -3

OpenStudy (gabylovesyou):

ok so r is -3

OpenStudy (gabylovesyou):

-3 * 4 = -12 -12^7 = -35,831,808

OpenStudy (johnweldon1993):

Right So now go back to the beginning \[\large a_n = a_1*r^{n-1}\] We now have a_1, we just found 'r'...we have all we need to solve for \(\large a_16\) \[\large a_{16} = a_1 * r^{16 - 1}\] \[\large a_{16} = 4*(-3)^{15}\]

OpenStudy (johnweldon1993):

You were still thinking about that 8th term we used in the beginning...that's trash now we dont need it now that we found 'r' :P Just remember...for these problems we need \(\large a_1\) and we need \(\large r\) and that's it Now that we have them...disregard everything else and only use those to find what you need

OpenStudy (gabylovesyou):

ok answer is B.. thank you..